OpenStudy (anonymous):

Bryan borrowed $1200 at 4 % simple interest rate for 4 months.He will repay the principal and interest in four equal payments.what is the amount of one payment?

OpenStudy (anonymous):

1200 times .04 (4%) equals 48 1200/4 (4 months) = 300 300+48 (monthly payment + interest) 348=one payment

OpenStudy (anonymous):

thanks. but the answer is 304

OpenStudy (anonymous):

I=Prt = 1200 *0.04*(4/12) = 16 ==> 1200+16 =1216 ==> payment = 1216/4=304

OpenStudy (anonymous):

thanks